OpenStudy (anonymous):

find the larger solution of 1/2x^2-20=92.5

OpenStudy (blacksteel):

Start by isolating x: 1/2x^2 - 20 = 92.5 x^2 - 40 = 185 x^2 = 225 Now solve the square for the solutions.

OpenStudy (anonymous):

is it 15 ?

OpenStudy (blacksteel):

Indeed it is! (The other solution is -15)

OpenStudy (campbell_st):

this simplifies to 1/2x^2 = 112.5 x^2 = 225 \[x=\pm 15\]

OpenStudy (anonymous):

thank you

OpenStudy (blacksteel):

You can always check by plugging your solution into the original equation: 1/2(15)^2 - 20 = 92.5 1/2(225) - 20 = 92.5 112.5 - 20 = 92.5 92.5 = 92.5

OpenStudy (campbell_st):

I would have thought that the solutions are equal in side...